A 65.0g sample of mg sulfide is found to contain 31.5g of mg. What is the percent by mass of sulfur in the compound?
1
Expert's answer
2020-10-19T14:26:03-0400

mass of sulfur = mass of MgS — mass of Mg

mass of sulfur = 65.0 g — 31.5 g = 33.5 g

"percent = \frac{mass of sulfur \times 100 }{mass of MgS}"

"persent = \frac{33.5 \times 100}{65.0} = 51.5"

Answer: 51.5 %